The length of the top of a table is 5m greater than the width. The area is 66m. Find the dimensions of the table
Answer by harpazo(655)

Width = x meters
Length = x + 5 meters
Area = 66 meters
Use Area = length times width
66 = (x+5)x
66 = x^2 + 5x
x^2 + 5x - 66 = 0
We now factor the quadratic equation to find x.
(x+11)(x-6) = 0
Note: Setting x+11 = 0 and solving for x yields
a negative answer. Since we are dealing
with dimension, reject x+11.

Solve x-6 = 0 for x.
In this case, x = 6.

Let's go back to the actual width and length
given above.
Width = x meters. We just found x to be 6. This
means the width is 6 meters.

Length = x + 5 meters. Letting x be 6, the length
turns out to be 6+5 meters or 11 meters.
Done.